# Towards Generalizing Inferences

### v1

@@ 5 Discussion and Conclusion
Our paper delves into the latest advancement in causal methods for generalizability and transportability, encapsulated by discussions at the ICERM workshop. It identifies three key areas of focus: (i) evaluating the robustness of foundational assumptions, (ii) extending the scope of causal analysis beyond traditional intention-to-treat effects, and (iii) incorporating machine learning to enhance the applicability of findings across various contexts.
The workshop also highlighted several challenges in bridging the research-to-practice gap. We provide examples of key areas requiring particular focus in the coming years:
@@
    Addressing Data Harmonization Challenges Most methods in the literature, including nearly all the work presented here, require data that is harmonized across trial and population. In other words, the data in the experiment and the target population must be measured in the same way and be encoded in the same way. Furthermore, the treatment administered in the experiment must also be administered in the same way in the target population. Constructing harmonized data poses a significant challenge and is not straightforward, especially given the current lack of coordination of measurement tools across trials and population data. A concerted effort is needed to design approaches for data harmonization, along with remedial strategies when harmonization is not feasible.
@@
    Addressing the Underrepresentation in Trials While many approaches discussed here focus on assessing violations of positivity and lack of representation, these approaches generally rely on posthoc adjustments and require strong assumptions. More work must be done to consider how to design future studies to provide information on target populations holistically, especially when the target population comprises of underrepresented groups.
@@
    Understanding the Underlying Causal Mechanism Inferring causal effects is not merely a statistical challenge; it heavily relies on underlying domain knowledge for the careful consideration of necessary assumptions. There is a need for a strong partnership between domain experts and methodologists to deepen the understanding of the underlying causal mechanisms and the plausibility of assumptions. This collaborative effort will contribute to developing robust and reliable approaches that align with real-world complexities.

### v2

@@ 5 Discussion and Conclusion
Our paper summarizes some of the recent advancements in causal methods for generalizability and transportability, as presented at the ICERM workshop. The workshop focused on three key areas of focus: (i) evaluating the robustness of foundational assumptions, (ii) extending the scope of causal analysis beyond traditional intention-to-treat effects, and (iii) incorporating machine learning to enhance the applicability of findings across various contexts. The discussion at the workshop also highlighted several challenges in bridging the research-to-practice gap as well as interesting future research directions. We provide a few examples of key discussion points from the workshop.
(1) Addressing Data Harmonization Challenges: Most methods in the literature, including nearly all the work presented here, require data that is harmonized across trial and population. In other words, the data in the experiment and the target population must be measured in the same way and be encoded in the same way. Furthermore, the treatment administered in the experiment must also be administered in the same way in the target population. Constructing harmonized data poses a significant challenge and is not straightforward, especially given the current lack of coordination of measurement tools across trials and population data. A concerted effort is needed to design approaches for data harmonization, along with remedial strategies when harmonization is not feasible.
(2) Addressing the Underrepresentation in Trials: While many approaches discussed here focus on assessing violations of positivity and lack of representation, these approaches generally rely on post hoc adjustments and require strong assumptions. More work must be done to consider how to design future studies to provide information on target populations holistically, especially when the target population comprises of underrepresented groups.
(3) Understanding the Underlying Causal Mechanism: Inferring causal effects is not merely a statistical challenge; it heavily relies on underlying domain knowledge for the careful consideration of necessary assumptions. There is a need for a strong partnership between domain experts and methodologists to deepen the understanding of the underlying causal mechanisms and the plausibility of assumptions. Furthermore, Tipton and Mamakos, (2023) highlighted the importance of moderators in recruiting a sufficiently heterogeneous experimental sample to generalize experimental results properly; however, a critical question remains of how researchers should determine what are important moderators a priori. Future work is needed to help establish systematic and data-driven approaches to help researchers determine salient characteristics in their target population that may moderate the treatment effect.
(4) Establishing Best Practice and Improving Interpretability: While recent methodological developments have introduced different tools for researchers to consider transporting or generalizing their causal effects, more work is needed to establish best practices and guidance for interpretability into the data analysis pipeline. For example, as discussed in Section 2, several of the proposed methods provided ways for researchers to integrate their substantive knowledge into assessing the validity of the underlying assumptions (i.e., Huang,, 2022’s benchmarking approach, as well as Zivich et al.,, 2023’s bounding method). While this provides an opportunity for researchers to leverage their substantive knowledge when considering the validity of the underlying identifying assumptions in more transparent and systematic ways, future work is needed to establish best practices and minimize potential misuse of these tools. For example, the pre-registration of moderators to benchmark could help minimize researcher bias when reporting sensitivity measures. Furthermore, many of the methodological approaches rely on complex statistical machinery that is often black-box to domain experts. Developing interpretable approaches by collaborating with domain experts would help improve the translational ability of the methods, saving time and effort while making the methods more applicable.
(5) Leveraging Connection to Recent ML Advancements: A separate but related field of research in machine learning literature, with significant overlap with causal inference, encompasses domain adaptation (Kouw and Loog,, 2018), distributionally robust optimization (Rahimian and Mehrotra,, 2019), and transfer learning (Pan and Yang,, 2009). These disciplines share the common challenges associated with generalizing or transporting inferences [^2] See Appendix for a brief note on the connections between causal inference and these domains. However, the lack of a shared vocabulary has thus far limited cross-disciplinary collaboration. A focused effort is needed to bridge the linguistic and research gaps between these areas, which would advance the development of new methodologies.
Appendix
Connections to Domain Adaptation, Distributionally Robust Optimization, Transfer Learning: In these areas, the training data used for fitting the model differs from the target dataset of interest, often unknown beforehand. In domain adaptation, the goal is to adapt a model trained on one domain to perform well on a different domain (Redko et al.,, 2019). This is often relevant in causal inference settings where the trial cohort may not perfectly represent the target population. Distributionally robust optimization aims to develop models that are robust to changes in the distribution of the data, ensuring that their performance remains consistent across different target populations (Duchi and Namkoong,, 2018; Rahimian and Mehrotra,, 2019; Staib and Jegelka,, 2019). Transfer learning focuses on transferring knowledge from a source task to a target task, even with different input or output spaces (Pan and Yang,, 2009; Zhuang et al.,, 2020). This approach is beneficial in scenarios where the target dataset is small or limited, allowing researchers to leverage insights from a related task to improve performance on the target task.
Workshop Talk Titles and Authors
@@
    Evaluating Ex Ante Counterfactual Predictions Using Ex Post Causal Inference (Gechter et al.,, 2018)
    Speaker: Cyrus Samii
@@
    Efficiently transporting average treatment effects using a sufficient subset of effect modifiers (Rudolph et al.,, 2023)
    Speaker: Kara Rudolph
@@
    Understanding effect heterogeneity in observational and randomized studies of causality (Vo et al.,, 2022)
    Speaker: Ivan Diaz
  4. 4.
    Generalizing trial evidence to target populations in non-nested designs: Applications to AIDS clinical trials (Li et al.,, 2022)
    Speaker: Ashley Buchanan
  5. 5.
    Extending Inferences to a Target Population Without Positivity (Zivich et al.,, 2023)
    Speaker: Paul Zivich
  6. 6.
    Generalizing and transporting inferences about the effects of the treatment assignment subject to non-adherence (Dahabreh et al.,, 2022)
    Speaker: Sarah Robertson
  7. 7.
    Who Are You Missing?: A Principled Approach to Characterizing the Underrepresented Population (Parikh et al.,, 2024)
    Speaker: Harsh Parikh
  8. 8.
    Using external data to address measurement error: a transportability perspective (ross2024leveraging)
    Speaker: Rachael Ross
  9. 9.
    Mitigating Bias in Treatment Effect Estimation: Strategies for Utilizing External Controls in Randomized Trials (Cheng et al.,, 2023)
    Speaker: Shu Yang
  10. 10.
    Improving Transportability of Randomized Clinical Trial Inference Using Robust Prediction Methods (Elliott et al.,, 2023)
    Speaker: Michael Elliott
  11. 11.
    Universal adaptability (Kim et al.,, 2022)
    Speaker: Frauke Kreuter
  12. 12.
    Sensitivity Analysis for Generalizing Experimental Results (Huang,, 2022)
    Speaker: Melody Huang
  13. 13.
    Beyond Generalization: Designing Randomized Experiments to Predict Treatment Effects (Tipton and Mamakos,, 2023)
    Speaker: Elizabeth Tipton
  14. 14.
    Extending Inferences from a Diverse Collection of Trials (Schnitzler and Kaizar,, 2023)
    Speaker: Eloise Kaizar
  15. 15.
    Transportability of Causal Effects in Principal Strata (Huling,, 2023)
    Speaker: Jared Huling